Periodic Inventory by Three Methods; Cost of Merchandise Sold The units of an item available for sale during the year were as follows: Jan. 1 Inventory 30 units @ $110 Mar. 10 Purchase 40 units @ $122 Aug. 30 Purchase 30 units @ $130 Dec. 12 Purchase 100 units @ $132 There are 60 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost and the cost of merchandise sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar. Cost of Merchandise Inventory and Cost of Merchandise Sold Inventory Method Merchandise Inventory Merchandise Sold First-in, first-out (FIFO) $ $ Last-in, first-out (LIFO) Weighted average cost
Answer #1
Unit Unit Cost Total Cost
Jan 1 30 110 3300
Mar 10 40 122 4880
Aug 30 30 130 3900
Dec 12 100 132 13200
Total 200 25280

Calculate following

FIFO LIFO Average cost
Cost of ending inventory 60*132 = 7920 (3300+30*122) = 6960 25280/200*60 = 7584
Cost of goods sold 25280-7920 = 17360 25280-6960 = 18320 25280-7584 = 17696